DVIDVI(L) DVIDVI(L) NAME dvidvi−selectsand/orre-arrangespagesinaTeXdvifile SYNOPSIS dvidvi[param]infileoutfile DESCRIPTION Copyright dvidvi1.0,Copyright(C)1988-91,RadicalEyeSoftware Introduction Thedvidviprogramconvertsadvifileintoanotherdvifile,withperhapscertainchanges. Parameters -fnpagenisfirstpageselected -lnpagenislastpageselected -nnselectatmostnpages.Noticethatnisthenumberofpagesselected,independentlyof thenumberofpagesactuallycontainedinasheet -i{n1..n2|n1}[,...] includepages(rangesallowed).Whenthisoptionisused,ONLYthespecifiedpagesareselected.However,wecanexcludefromthesepageswiththe option-x -x{n1..n2|n1}[,...] excludepages(rangesallowed) -qworkinquietmode,thatisdonotprintinthescreenmessagesofhowtheworkisbeing done. -rreversetheorderofthepages. Thepagenumbersfortheaboveoptions-f-l-iand-xcanbespecifiedindifferentways. 1)Ifanumbernisgiven,itisinterpretedasthen’thpagefromthebeginningofthe.dvifile.Of course,thisnumberisindependentofthepagenumberassignedbyTeX. 2)TeXpagenumbersarethosewhoareactuallywritteninthepage;thesepagenumberscanbe modified,forexample,byusingtheTeXcommands\pagenumbering,\setcounter{page}{n},and\addtocounter{page}{n}.ATeXpagenumbercanbespecified byprecedingthenumbernwiththecharacter@.Thus,ifyouspecify-f@25-l@30you selectthepagesbetween25and30,thesenumbersbeingthoseassignedbyTeX. 3)However,severalpagescanhavethesameTeXpagenumberina.dvifile.Forexample,the introductorypagesinabookarenumberedi,ii,andsoonuntilthefirstchapterbegins andthen,thepagesarenumbered1,2,etc.Inthiscase,thepagesnumberediand1in the.dvifilehavethesameTeXpagenumber.Ifyouwanttoselectforexamplethesecondoccurrenceofthepagenumbered1,youcanspecifyapagenumberas(@2)1.Thus @1isequivalentto(@1)1.Forexample,ifyouspecify-f(@2)1-l(@2)10youselectthe pagesbetween1and10ofthefirstchapter,nottheintroductorypagesbetweeniandx. Thereisanotherparameterthattellsdvidvihowyouwanttochangepagelayoutandspecifications.Thisisthe-mparameter. *Thenumberprecedingthecolonisthemodulovalue.Everythingwillbedoneinchunksof pagesthisbig.Ifthereisnocolon,thanthedefaultvalueisassumedtobeone.The lastchunkofpagesispaddedwithasmanyblankpagesasnecessary. *Followingthecolonisacomma-separatedlistofpagenumbers.Thesepagenumbersarewith respecttothecurrentchunkofpages,andmustlieintherangezerotothemodulovalue lessone.Ifanegativesignprecedesthenumber,thenthepageistakenfromthemirror chunk;iftherearemchunks,thenthemirrorchunkofchunknisthechunknumbered m-n-1.Putsimply,itisthechunknumberedthesame,onlyfromtheend.Thiscanbe usedtoreversepages.Ifnonumberisgiven,thepagenumberdefaultsto1. *Followingeachpagenumberisanoptionaloffsetvalueinparenthesis,whichconsistsofapair ofcomma-separateddimensions.Eachdimensionisadecimalnumberwithanoptional unitofmeasure.Thedefaultunitofmeasureisinches,orthelastunitofmeasureused. March1994 1 DVIDVI(L) DVIDVI(L) Allunitsareintruedimensions.AllowableunitsofmeasurearethesamethatTeX allows:in,mm,cm,pt,pc,dd,andcc. EXAMPLES -m-Reversestheorderofthepages.Thistime,boththemoduloandthepagenumberare defaulted. -m2:0Selectsthefirst,third,fifth,etc.pagesfromthefile.Printthisoneafterprintingthe next,takingthepaperoutofthefeedtrayandreinsertingitintothepaper feed. -m2:-1Selectsthesecond,fourth,etc.pages,andwritestheminreverseorder. -m4:-1,2(4.25in,0in) -m4:-3,0(4.25in,0in) Usefulforprintingalittlebooklet,fourpagestoasheet,double-sided,for staplinginthemiddle.Printthefirstone,putthestackbackintothe printerupsidedown,andprintthesecond.The‘in’specificationsare superfluous. -m,(1pt,1) Scareyoursystemadministrator!Actually,thingsaresoblurrywiththis option,youmaywanttosendenemieslettersprintedlikethis.*Long*letters. -m4:0(5.5in,4.25),3(0,4.25) -m4:1(0in,4.25),2(5.5,4.25) Printafour-pagecardononesheet.Printthefirst,rotatethepaper180 degreesandfeeditagain.(PostScriptpeoplecandofunnytrickswith PostScriptsothisisn’tnecessary.) March1994 2